Three independent samples of ten digits 0,1,2....,9 are taken.Determine the probability that the same digit occurs more than once in the three samples?
Answer by sudhanshu_kmr(1152)


probability that no repetition of digit any digit i.e P = 10/10* 9/10 * 8/10

= 72/100
now, probability that same digit occurs more than once i.e P' = 1 - P
= 1 -72/100
= 28/100